OpenStudy (anonymous):

Final question guys! Need this for a good grade please help me out. I need help finding this equation/ An investor bought some municipal bonds yielding 5% annually and some certificates of deposit yielding 7% annually. If her total investment amounts to $20,000 and her annual return is $1160, how much money is invested in bonds and how much in certificates of deposit?

OpenStudy (anonymous):

Plato much?

OpenStudy (anonymous):

.05b + .07c = 1160 b + c = 20000

OpenStudy (anonymous):

Excuse me?

OpenStudy (anonymous):

Ty Mramirez I am going to try and do this :) IOU

OpenStudy (anonymous):

No problem, let me know if you get stuck

OpenStudy (anonymous):

Well I am wondering How will I solve for b and c?? I am thinking about this and I havent done tthese types of problems 5 years.

OpenStudy (anonymous):

You can use either substitution or elimination for this 2x2 linear system.

OpenStudy (anonymous):

I would probably use substitution because you can get a variable alone by itself

OpenStudy (anonymous):

b + c = 20000 -c -c b=20000 - c insert that into the other equation

OpenStudy (anonymous):

.05(20000 - c) + .07c = 1160

OpenStudy (anonymous):

Ok that clears things up a ton more.

OpenStudy (anonymous):

How do i get the 2cs combined or dont i?

OpenStudy (anonymous):

.05(20000 - c) + .07c = 1160 1000-.05c + .07c = 1160 1000 + .02c = 1160 -1000 -1000 .02c=160 /.02 /.02 c=8000

OpenStudy (anonymous):

Insert c into the other equation to solve for b

OpenStudy (anonymous):

b + 8000 = 20000

OpenStudy (anonymous):

I found that equation just as you submited. You are an awesome help. I found a useful link http://www.purplemath.com/modules/systlin4.htm What i was going off of as well. Ty again mramirez :)

OpenStudy (anonymous):

No worries, have a great day

OpenStudy (anonymous):

You too !